# AGN - host galaxy photometric decomposition using a fast, accurate and precise deep learning approach

### v1

@@ 2.2 Mock JWST images
Figure 2: Example mock JWST F150W images with varying levels of AGN contribution. The images have been generated to mimic JWST observations, and include realistic JWST noise and background. Each row corresponds to a different galaxy with no AGN contribution in the left panel and increasing AGN contributions in the rest of the panels. We show four example galaxies with different magnitudes, from the brightest (top) to the faintest (bottom). Images correspond to a physical size of 50 kpc and are displayed with an inverse arcsinh scaling.
The IllustrisTNG project (Nelson et al., 2019; Pillepich et al., 2018; Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) is a series of cosmological hydrodynamical simulations of galaxy formation and evolution, with three different runs that differ in volume and resolution. These runs are TNG50, TNG100, and TNG300, with comoving length sizes of 50, 100, and 300 Mpc $h^{-1}$, respectively. The initial conditions for all runs are drawn from Plank results (Planck Collaboration et al., 2016). For this work, we used TNG100 (better resolution than TNG300 but still with a big enough volume to contain a sufficiently large number of galaxies), which contains $1820^{3}$ dark matter (DM) particles with a mass resolution of $M_{\mathrm{DM,\ res}}=7.5\times 10^{6}M_{\odot}$. In comparison, the baryonic particle resolution of TNG100 is $M_{\mathrm{baryon\,res}}=1.4\times 10^{6}$ M. We refer the reader to Pillepich et al. (2018) for more details on IllustrisTNG.
In this work, we selected galaxies from six simulation snapshots (numbers 67, 64, 61, 58, 55 and 52), which correspond to redshifts from $z=0.5$ out to $0.92$. The time step between each snapshot is roughly $\sim$480 Myr over this redshift interval. We selected all galaxies from these six snapshots with stellar mass $M_{*}>10^{9}M_{\odot}$ to ensure that most galaxies have a sufficient number of stellar particles (hence reasonably well resolved), with the lowest mass galaxies in TNG100 ($M_{*}=10^{9}M_{\odot}$) consisting of $714$ particles. This selection in mass and redshift resulted in a total of $\sim 120000$ galaxies. Every galaxy in IllustrisTNG has a complete merger history available from applying the SUBLINK algorithm on baryon-based structures (Rodriguez-Gomez et al., 2015). We used these merger trees to identify major mergers and non-merger galaxies. Specifically, major mergers are defined as galaxies with stellar mass ratios ¿ 1:4 and that will either have a merger event in the following 0.8 Gyr (pre-mergers) or had a merger event in the last 0.3 Gyr (post-mergers). Consequently, galaxies that do not satisfy these conditions are considered as non-mergers.
@@
In the final step, we followed the same approach as in Margalef-Bentabol et al. (2024). To obtain cutouts of the real JWST sky, we first generated a catalogue of sources that we wanted to avoid within the central region of the cutouts. Starting from the ‘Farmer’ version of the COSMOS2020 catalogue (Weaver et al., 2022), we set the flag flag_combined equal to 0 to select areas that are not affected by bright stars or large artefacts, as recommended by the COSMOS team. We then made use of the star/galaxy separation provided in the Le Phare photo-$z$, selecting sources with the star/galaxy flag lp_type equal to 0 (galaxy) or 2 (X-ray source). Finally, we restricted our selection only to $z<3$ sources in the COSMOS-Web field. The final catalogue contains relatively bright sources, with average magnitudes being: HSC-$i=25.6$ mag, UVISTA-$H=24.7$ mag. Then, we generated random sky coordinates such that there are no catalogued bright sources within a circular radius of $6.5\arcsec$. This radius corresponds to the estimated source density of the area from which we extracted the cutouts. These criteria ensure that there are no bright sources in the centre of the cutouts, where the synthetic galaxies will be injected, but still allow for faint background galaxies. The random coordinates were then used as the centres for the sky cutouts, in which to inject the simulated galaxies, after performing sanity checks to ensure that no artefacts, stellar spikes, or bad or saturated pixels were present. For different snapshots of the simulation (and therefore, at different redshifts), a physical size of 50 kpc corresponds to a different image size in pixels (ranging from 210 to 270 pixels). To have a uniform dataset to train our deep learning model, we resized all images to 256 pixels. This transformation effectively changes the image pixel resolution depending on the corresponding snapshot/redshift, from 0.025″/ pixel at the highest redshift to 0.032″/ pixel at the lowest redshift. In Section 2.2, where we injected the observed PSF models, the pixel resolution of the PSF model was adjusted accordingly, to exactly match the pixel resolution of the mock images of simulated galaxies at different redshifts.
@@ 3 Methods
In this section, we first describe the construction of mock host galaxy images with different injected levels of AGN contribution. Then we introduce the two methods (our method based on DL and GALFIT based on 2D surface brightness fitting) used to estimate the AGN contribution.
@@ 3.1 Mock AGN injection
To simulate images of galaxies with AGNs, we injected a central point source into the host galaxy image. The observed JWST PSF models as described in Section 2.1 were used as the central point source. In order to create different AGN fractions, the relative brightness of the PSF was adjusted before injecting it into the mock JWST images described in Section 2.2.
The AGN fractions were chosen to range from 0.0 to 0.95 in increments of 0.05, for a total of 20 different AGN fractions. For any given AGN fraction $f_{AGN}$, the injected image was created as follows. First, the PSF image was transformed into the correct pixel scale for each redshift then, the flux of the PSF image was measured within a 3″aperture [^3] A $3$aperture corresponds to different sizes in pixels for different redshift, due to having resized all images to the same number of pixels. using the aperture_photometry function of the photutils package (Bradley et al., 2024). This aperture size corresponds to sizes between $18.6$ kpc and $23.9$ kpc in our redshift range. The PSF image was normalised using this flux, so it could be easily adjusted later. The flux of the host galaxy image, $F_{host}$, was measured in the same way. The AGN fraction (i.e. contribution to the total light) is then defined as
@@
where $F_{AGN}$ is the flux of the AGN. Afterwards, the normalised PSF image was multiplied by $F_{AGN}$, which can be derived from Equation 1:
@@
Finally, the transformed PSF image was added to the host galaxy image. For each galaxy, five different images were created with five different AGN fractions, chosen randomly from the 20 possible discrete AGN fractions. Some example images of these simulated galaxies (from bright to faint) with varying levels of AGN contributions can be seen in Fig. 2. Visually, it is clear that it can be very difficult to discern the presence of the host galaxy when $f_{AGN}\geq 0.8$, particularly for fainter and higher-redshift objects. Later we will examine how well we can quantify $f_{AGN}$ as a function of redshift, magnitude, signal-to-noise, etc.
Figure 4: Distribution of injected AGN fraction (as defined in Equation 1) in the training (blue) and test datasets (orange). The distributions are mostly uniform for the training and test datasets for the AGN fraction.
@@ 3.2 Deep learning CNN
Zoobot (Walmsley et al., 2023) is a Python package used to measure detailed morphologies of galaxies (such as spiral arms, bars and bulges) using DL based on the idea of successive layers of learned representations. Zoobot includes convolutional neural networks (CNNs, Fukushima, 1988; LeCun et al., 2015) and vision transformer models (Dosovitskiy et al., 2021; Dehghani et al., 2023). These models are pre-trained on many millions of labelled galaxies, derived from the visual classifications of the Galaxy Zoo project(Lintott et al., 2008) on real images of galaxies selected from surveys such as the Sloan Digital Sky Survey (SDSS), Hyper Suprime-Cam (HSC) and Hubble (Willett et al., 2013, 2017; Simmons et al., 2017; Walmsley et al., 2022a, b). The models are designed to be easily adaptable to new tasks (classification or regression tasks) and galaxy surveys with a minimal amount of new labelled data. Some of the available models in Zoobot belong to the family of ConvNeXts (Liu et al., 2022). They are pure convolutional models constructed by optimising the ResNet (He et al., 2015) architecture to bear resemblance with vision transformers (Vaswani et al., 2017; Liu et al., 2021), in which the design choices such as the use of the Gaussian Error Linear Unit (GELU) activation function or inverted bottleneck CNN blocks (which are a specialised type of residual block more computationally efficient than normal residual blocks) are proven to improve the performance of a purely CNN model and can compete with transformer in terms of accuracy and scalability. For this work, we chose a ConvNeXt-Base architecture, which was pre-trained on the Galaxy Zoo dataset of over 820k images and 100 million volunteer votes to morphological questions. In order to perform a regression task (as is the case in this study), we added a linear head with a sigmoid function (to restrict the output to be between 0 and 1) and used a mean square error loss function to train the network. A diagram of the ConvNeXt-Base network can be seen in Fig. 3. For this work, we retrained the last two blocks of the network and the linear head, while the rest of the network parameters were frozen to keep the optimal values found for the pre-trained data from Galaxy Zoo.
@@
We split our sample of mock JWST images in training and test sets, with a 90/10 split. The distribution of the injected AGN fraction on the training and test datasets can be seen in Fig. 4. The training dataset was used for training Zoobot. A validation set of 10% of the training data was used to monitor the performance during training. The test set was used to determine the performance of the final model on data that had never been seen before by the algorithms. To ensure the test set cannot be learned by simply interpolating from the training set, the galaxies were split into the train, validation and test sets in a way that the five iterations of the same galaxy (with different injected AGN fractions) were only used in one of the splits. Furthermore, we randomly selected a subset of $2400$ galaxies from the test set on which to directly compare the results from the DL methodology and surface brightness fitting with GALFIT. This subset was constructed in a way that for each snapshot we randomly selected 400 galaxies with a uniform distribution of $f_{AGN}$.
@@ 3.3 GALFIT
Figure 5: Sérsic + PSF decomposition of four example galaxies (with AGN fraction varying from high to low from top to bottom) on which we performed GALFIT. Images of the original galaxy, the model (Sérsic + PSF), and the residual (original--model) are shown from left to right. Images correspond to a physical size of 50 kpc, and are displayed with an inverse arcsinh scaling.
@@
We ran GALFIT on the subset of 2400 galaxies from the test set. We first ran Sextractor (Bertin & Arnouts, 1996) on all the test images to determine the central positions in pixels, x and y, the total magnitude of the galaxy, axis ratio (q) and effective radius ($R_{e}$). Sextractor also produces a segmentation map in which all galaxies are identified and can be used to either mask neighbouring galaxies or fit them simultaneously with the central galaxy. In order to perform $\chi^{2}$-minimisation, GALFIT requires a sigma map. This sigma map was constructed by adding the Poisson noise contribution from the simulated galaxy (after injecting the AGN contribution) to the error map provided by the JWST data.
@@
In some cases, GALFIT did not converge and did not produce a fit at all. In other cases, even if GALFIT produced an output, it was clearly not a good fit for the surface brightness light. We only selected galaxies for which GALFIT produced a good fit, that is, it has a reduced $\chi^{2}<5$ and no non-physical parameters (for example effective radius smaller than 0.5 pixels, or larger than the size of the image stamp, $q<0.1$, and $n<0.5$ or $n>10$). In Fig. 5, we present good GALFIT fits of example galaxies containing varying amounts of injected AGN fractions (from insignificant to dominant AGN contributions), with the original image, the model image (Sérsic + PSF) and the residual image (original - model) shown in different columns. After the best fit had been found, we created images of each model separately (Sérsic and PSF) using the parameters from the best fit. We then calculated the flux within an aperture of 3″in the Sérsic model, the PSF model and the original galaxy image.
Finally, we calculated the AGN fraction derived from running GALFIT in two slightly different ways. In the first method, we set the derived AGN fraction to be equal to the ratio of the aperture flux from the PSF component to the aperture flux of the original galaxy image (which corresponds to the total flux of the galaxy within a 3″aperture). In the second method, we used as total flux (within a 3″aperture) the sum of the aperture fluxes of the PSF and the Sérsic model. We adopted these two different approaches to calculating the total flux of the galaxy in order to understand if there is any systematic bias in the Sérsic model, which could also impact the flux of the PSF model.
@@ 4 Results
In this section, we first present results from our DL-based method Zoobot. We analyze in detail how key metrics such as the root mean square error (RMSE), relative absolute error (RAE) and outlier fraction vary as a function of the injected AGN fraction, redshift, signal-to-noise (S/N), and magnitude. Then we compare these results with the performance obtained from running GALFIT on the same test set.
@@ 4.1 Zoobot model performance
Figure 6: Comparison between the real injected AGN fraction and the AGN fraction obtained from the Zoobot model on the subset of 2400 galaxies across the whole redshift range ($10.5<z<$ 1). The comparison shows a mean difference between the two quantities ($\Delta f_{AGN}=f_{AGN}$ [Injected] $-f_{AGN}$ [Zoobot]) of $-0.0006$ and an overall RMSE = 0.027. The solid diagonal line is the 1:1 line. The top plot shows the mean difference and its dispersion as a function of the injected AGN fraction. The colour bar indicates the number of sources in each 2D bin.
@@
Zoobot $0.027\pm 0.005$ $0.086\pm 0.004$ $-0.0006\pm 0.0005$ $0.078\pm 0.006$ $0.063\pm 0.005$
ZoobotM $0.019\pm 0.004$ $0.085\pm 0.009$ $-0.0021\pm 0.0008$ $0.07\pm 0.01$ $0.06\pm 0.01$
ZoobotNM $0.029\pm 0.005$ $0.087\pm 0.005$ $-0.0001\pm 0.0007$ $0.081\pm 0.007$ $0.064\pm 0.006$
GALFIT ($f_{host}$ from original image) $0.090\pm 0.008$ $0.19\pm 0.01$ $-0.018\pm 0.002$ $0.19\pm 0.01$ $0.130\pm 0.009$
GALFITM ($f_{host}$ from original image) $0.090\pm 0.009$ $0.27\pm 0.05$ $-0.024\pm 0.004$ $0.25\pm 0.03$ $0.18\pm 0.02$
GALFITNM ($f_{host}$ from original image) $0.090\pm 0.008$ $0.17\pm 0.01$ $-0.016\pm 0.002$ $0.16\pm 0.01$ $0.113\pm 0.009$
GALFIT ($f_{host}$ from model) $0.141\pm 0.009$ $0.38\pm 0.02$ $-0.072\pm 0.003$ $0.37\pm 0.02$ $0.27\pm 0.01$
GALFITM ($f_{host}$ from model) $0.135\pm 0.008$ $0.44\pm 0.07$ $-0.082\pm 0.005$ $0.47\pm 0.04$ $0.33\pm 0.03$
GALFITNM ($f_{host}$ from model) $0.142\pm 0.008$ $0.36\pm 0.02$ $-0.069\pm 0.003$ $0.34\pm 0.02$ $0.25\pm 0.01$
@@
Here we analyse the performance of the Zoobot model on the subset of 2400 galaxies from the test set. We first calculate the root mean square error (RMSE),
@@
which measures the average difference between the predicted values from a model (e.g. Zoobot or GALFIT) and the actual injected values. In Fig. 6 we show the predicted values of $f_{AGN}$, obtained from Zoobot, versus the real $f_{AGN}$ for all redshifts, which shows a very tight correlation across the whole dynamic range. Quantitatively, we found an overall value of $\text{RMSE}=0.027$ for the Zoobot model, demonstrating very good recovery (both in terms of accuracy and precision) of the injected real AGN fractions. In addition, the top panel of Fig. 6 shows the mean and the dispersion of the difference between the real and predicted values ($\Delta f_{AGN}=f_{AGN}$ [injected] $-f_{AGN}$ [Zoobot]) as a function of the injected AGN fraction, with a value for the mean of $\langle{\Delta f_{AGN}}\rangle=-0.0006$, and dispersion of $\sigma(\Delta f_{AGN})=0.027$. This shows that there is very little bias from the Zoobot predictions to the actual values, and it is independent of $f_{AGN}$, as the difference between the real and predicted value over the whole range of $f_{AGN}$ is always below -0.006. Moreover, the overall dispersion is very small, particularly for the high and low injected AGN fractions ($f_{AGN}\gtrsim 0.8$ and $f_{AGN}\lesssim 0.2$). To check if the performance depends on galaxy structural properties, we further explore whether the Sérsic index impacts how well Zoobot can predict the AGN fraction (see Fig. 12 in Appendix A) and find that for galaxies with $n<1$ the RMSE increases by 30%, to RMSE$=0.035$ and that high Sérsic indexes ($n>1$) do not lead to worse predictions (RMSE $=0.020$).
@@
Figure 7: RMSE as a function of the injected AGN fraction (top left panel), redshift (top right), signal-to-noise (bottom left) and magnitude (bottom right). The purple lines correspond to the results from Zoobot and the green lines from GALFIT. The solid lines correspond to the whole sample while the dashed and dotted lines correspond to the mergers and non-merger galaxies, respectively. The error bars show the 95% interval from bootstrapping the RMSE value. The performance of Zoobot is around a factor of 3 - 10 better than GALFIT. In general, highly disturbed or irregular galaxies such as mergers are more difficult for both methods. The shaded region on the top left panel represents the fractional variation (standard deviation divided by the mean) of the PSF, considering the spatial and temporal variations, showing how Zoobot is able to achieve smaller errors than the errors from the variation of the PSF.
@@
Figure 8: Average relative absolute error in bins of injected $f_{AGN}$ (top left panel), redshift (top right panel), signal-to-noise (bottom left panel) and magnitude (bottom right), from the Zoobot predictions. The error bars correspond to the standard deviation of the data points in each bin. The purple lines correspond to the results from Zoobot and the green, from GALFIT. The solid lines correspond to the whole sample while the dashed and dotted lines correspond to the merger and non-merger galaxies, respectively.

### v2

@@ 2.3 Mock JWST images
Figure 2: Example mock JWST/NIRCAM F150W images with varying levels of AGN contribution. The images have been generated to mimic JWST observations, and include realistic JWST noise and background. Each row corresponds to a different galaxy with no AGN contribution in the left panel and increasing AGN contributions in the rest of the panels. We show four example galaxies with different magnitudes, from the brightest (top) to the faintest (bottom). Images are 3.84″across and are displayed with an inverse arcsinh scaling.
The IllustrisTNG project (Nelson et al., 2019; Pillepich et al., 2018; Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) is a series of cosmological hydrodynamical simulations of galaxy formation and evolution, with three different runs that differ in volume and resolution. These runs are TNG50, TNG100, and TNG300, with comoving length sizes of 50, 100, and 300 Mpc $h^{-1}$, respectively. The initial conditions for all runs are drawn from the Plank results (Planck Collaboration et al., 2016). For this work, we used TNG100 (better resolution than TNG300 but still with a big enough volume to contain a sufficiently large number of galaxies), which contains $1820^{3}$ dark matter (DM) particles with a mass resolution of $M_{\mathrm{DM,\ res}}=7.5\times 10^{6}M_{\odot}$. In comparison, the baryonic particle resolution of TNG100 is $M_{\mathrm{baryon\,res}}=1.4\times 10^{6}$ M. We refer the reader to Pillepich et al. (2018) for more details on IllustrisTNG.
In this work, we selected galaxies from fourteen simulation snapshots (numbers 67, 64, 61, 58, 55, 52, 49, 46, 43, 40, 37, 33, 29 and 25), which correspond to redshifts from $z=0.5$ out to $3$. The time step between each snapshot is roughly $\sim$480 Myr over this redshift interval. We selected all galaxies from these snapshots with stellar mass $M_{*}>10^{9}M_{\odot}$ to ensure that most galaxies have a sufficient number of stellar particles (hence reasonably well resolved), with the lowest mass galaxies in TNG100 ($M_{*}=10^{9}M_{\odot}$) consisting of $714$ particles. We randomly choose a total of $\sim 12000$ galaxies from this selection in mass and redshift to create our training sample. Every galaxy in IllustrisTNG has a complete merger history available from applying the SUBLINK algorithm on baryon-based structures (Rodriguez-Gomez et al., 2015). We used these merger trees to identify major mergers and non-merger galaxies. Specifically, major mergers are defined as galaxies with stellar mass ratios ¿ 1:4 and that will either have a merger event in the following 0.8 Gyr (pre-mergers) or had a merger event in the last 0.3 Gyr (post-mergers). Consequently, galaxies that do not satisfy these conditions are considered as non-mergers.
@@
In the final step, we followed the same approach as in Margalef-Bentabol et al. (2024). To obtain cutouts of the real JWST sky, we first generated a catalogue of sources that we wanted to avoid within the central region of the cutouts. Starting from the ‘Farmer’ version of the COSMOS2020 catalogue (Weaver et al., 2022), we set the flag flag_combined equal to 0 to select areas that are not affected by bright stars or large artefacts, as recommended by the COSMOS team. We then made use of the star/galaxy separation provided in the Le Phare photo-$z$, selecting sources with the star/galaxy flag lp_type equal to 0 (galaxy) or 2 (X-ray source). Finally, we restricted our selection only to $z<3$ sources in the COSMOS-Web field. The final catalogue contains relatively bright sources, with average magnitudes being: HSC-$i=25.6$ mag, UVISTA-$H=24.7$ mag. Then, we generated random sky coordinates such that there are no catalogued bright sources within a circular radius of $6.5\arcsec$. This radius corresponds to the estimated source density of the area from which we extracted the cutouts. These criteria ensure that there are no bright sources in the centre of the cutouts, where the synthetic galaxies will be injected, but still allow for faint background galaxies. The random coordinates were then used as the centres for the sky cutouts, in which to inject the simulated galaxies, after performing sanity checks to ensure that no artefacts, stellar spikes, or bad or saturated pixels were present.
To have a uniform dataset to train our DL model, and to speed up computation, we cut all images to $128\times 128$ pixels, without changing the pixel scale of 0.03″/pixel. In other words, the images are 3.84″across, corresponding to roughly 23 - 33 kpc in physical scale over the redshift range of our sample (i.e. $30.5<z<$ 3).
@@ 3 Methods
In this section, we first describe the construction of the mock JWST/NIRCAM host galaxy images with different injected levels of AGN contribution. Then we introduce the two methods (our method based on DL and GALFIT based on 2D surface brightness fitting) used to estimate the AGN contribution fractions.
@@ 3.1 Mock AGN injection
To simulate images of galaxies with AGN, we injected a central point source into the host galaxy image. The observed JWST PSF models as described in Section 2.1 were used as the central point source. In order to create different AGN contribution fractions, the relative brightness of the PSF was adjusted before injecting it into the mock JWST images described in Section 2.3.
The AGN contribution fractions were chosen to range from 0.0 to 0.95 in increments of 0.05, for a total of 20 different levels. For any given AGN contribution fraction $f_{AGN}$, the injected image was created as follows. First, the PSF image was transformed into a pixel scale of 0.03″/pix, same as the simulated images. Then, the flux of the PSF image was measured within a 2″aperture using the aperture_photometry function of the photutils package (Bradley et al., 2024). This aperture size corresponds to sizes between $13$ kpc and $18$ kpc in our redshift range. The PSF image was normalised using this flux, so it could be easily adjusted later. The flux of the host galaxy image, $F_{host}$, was measured in the same way. The AGN contribution fraction (i.e. contribution to the total light) is then defined as
@@
where $F_{AGN}$ is the flux of the AGN. Afterwards, the normalised PSF image was multiplied by $F_{AGN}$, which can be derived from Eq. 7:
@@
Finally, the scaled PSF image was added to the host galaxy image.
For each galaxy, five different images were created with five different AGN contribution fractions, chosen randomly from the 20 possible discrete AGN contribution fractions. Example images of these simulated galaxies (from bright to faint) with varying levels of AGN contribution can be seen in Fig. 2. Visually, it is clear that it can be very difficult to discern the presence of the host galaxy when the AGN contribution fraction is large, particularly for fainter and higher-redshift objects. Later we will examine how well we can recover $f_{AGN}$ as a function of the intrinsic AGN contribution fraction, redshift, galaxy size, S/N, etc, using a set of common performance metrics.
Figure 4: Distributions of the injected AGN contribution fraction (as defined in Eq. 7) in the training (blue) and test datasets (orange). The distributions are mostly uniform for both the training and test datasets.
@@ 3.2 Deep learning CNN
Zoobot (Walmsley et al., 2023) is a Python package used to measure detailed morphologies of galaxies (such as spiral arms, bars and bulges) using DL based on the idea of successive layers of learned representations. Zoobot includes convolutional neural networks (CNNs, Fukushima, 1988; LeCun et al., 2015) and vision transformer models (Dosovitskiy et al., 2021; Dehghani et al., 2023). These models are pre-trained on many millions of labelled galaxies, derived from the visual classifications of the Galaxy Zoo project(Lintott et al., 2008) on real images of galaxies selected from surveys such as the Sloan Digital Sky Survey (SDSS), Hyper Suprime-Cam (HSC) and Hubble (Willett et al., 2013, 2017; Simmons et al., 2017; Walmsley et al., 2022a, b). The models are designed to be easily adaptable to new tasks (classification or regression tasks) and galaxy surveys with a minimal amount of new labelled data.
Some of the available models in Zoobot belong to the family of ConvNeXts (Liu et al., 2022). They are pure convolutional models constructed by optimising the ResNet (He et al., 2015) architecture to bear resemblance with vision transformers (Vaswani et al., 2017; Liu et al., 2021), in which the design choices such as the use of the Gaussian Error Linear Unit (GELU) activation function or inverted bottleneck CNN blocks (which are a specialised type of residual block more computationally efficient than normal residual blocks) are proven to improve the performance of a purely CNN model and can compete with transformer models in terms of accuracy and scalability.
For this work, we chose a ConvNeXt-Base architecture, which was pre-trained on the Galaxy Zoo dataset of over 820k images and 100 million volunteer votes to morphological questions. In order to perform a regression task (as is the case in this study), we added a linear head with a sigmoid function (to restrict the output to be between 0 and 1) and used a mean square error loss function to train the network. A diagram of the ConvNeXt-Base network can be seen in Fig. 3. For this work, we retrained the last two blocks of the network and the linear head, while the rest of the network parameters were frozen to keep the optimal values found for the pre-trained data from Galaxy Zoo.
@@
We split our sample of mock JWST images in training and test sets, with a 90/10 split. The distribution of the injected AGN contribution fraction in the training and test datasets can be seen in Fig. 4. The training dataset was used for training Zoobot. A validation set of 10% of the training data was used to monitor the performance during training. The test set was used to determine the performance of the final model on data that had never been seen before by the algorithms. To ensure the test set cannot be learned by simply interpolating from the training set, the galaxies were split into the train, validation and test sets in a way that the five iterations of the same galaxy (with different injected AGN contribution fractions) were only used in one of the splits. Furthermore, we randomly selected a subset of $4800$ galaxies from the test set on which to directly compare the results from the DL methodology and surface brightness fitting with GALFIT. This subset was constructed in a way that for each snapshot we randomly selected 400 galaxies with a uniform distribution of $f_{AGN}$.
@@ 3.3 GALFIT
Figure 5: Sérsic + PSF decomposition of four example galaxies (with AGN contributon fraction varying from high to low from top to bottom) on which we performed GALFIT. Images of the original galaxy, the model (Sérsic + PSF), and the residual (original--model) are shown from left to right. Images are 3.84″across, and are displayed with an inverse arcsinh scaling.
@@
We ran GALFIT on the subset of 4800 galaxies from the test set. We first ran Sextractor (Bertin & Arnouts, 1996) on all the test images to determine the central positions in pixels, x and y, the total magnitude of the galaxy, axis ratio (q) and effective radius ($R_{e}$). Sextractor also produces a segmentation map in which all galaxies are identified and can be used to either mask neighbouring galaxies or fit them simultaneously with the central galaxy. In order to perform $\chi^{2}$-minimisation, GALFIT requires a sigma map. This sigma map was constructed by adding the Poisson noise contribution from the simulated galaxy (after injecting the AGN contribution) to the error map provided by the JWST data.
@@
In some cases, GALFIT did not converge and did not produce a fit at all. In other cases, even if GALFIT produced an output, it was clearly not a good fit for the surface brightness light. We only selected galaxies for which GALFIT produced a good fit, that is, it has a reduced $\chi^{2}<5$ and no non-physical parameters (for example effective radius smaller than 0.5 pixels, or larger than the size of the image stamp, $q<0.1$, and $n<0.5$ or $n>10$). In Fig. 5, we present good GALFIT fits of example galaxies containing varying amounts of injected AGN contribution fractions (from insignificant to dominant AGN contributions), with the original image, the model image (Sérsic + PSF) and the residual image (original - model) shown in different columns. After the best fit had been found, we created images of each model separately (Sérsic and PSF) using the parameters from the best fit. We then calculated the flux within an aperture of 2″in the Sérsic model, the PSF model and the original galaxy image.
Finally, we calculated the AGN contribution fraction derived from running GALFIT in two slightly different ways. In the first method, we set the derived AGN contribution fraction to be equal to the ratio of the aperture flux from the PSF component to the aperture flux of the original galaxy image (which corresponds to the total flux of the galaxy within a 2″aperture). In the second method, we used as total flux (within a 2″aperture) the sum of the aperture fluxes of the PSF and the Sérsic model. We adopted these two different approaches to calculating the total flux of the galaxy in order to understand if there is any systematic bias in the Sérsic model, which could also impact the flux of the PSF model.
@@ 4 Results
In this section, we first present results from our DL-based method Zoobot. We analyse in detail how key performance metrics such as the root mean square error (RMSE), relative absolute error (RAE) and outlier fraction vary as a function of the injected AGN contribution fraction, redshift, S/N, and galaxy size. Then, we compare these results with the performance obtained from running GALFIT on the same test set. Finally, we apply our DL-based method to the real JWST images of the stellar mass complete galaxy sample and compare AGN identified using our method with AGN selected in the X-ray and MIR.
@@ 4.1 Zoobot model performance
Figure 6: Comparison between the real injected AGN contribution fraction and the AGN contribution fraction obtained from the Zoobot model on the subset of 4800 galaxies across the whole redshift range ($30.5<z<$ 3). The comparison shows a mean difference between the two quantities ($\Delta f_{AGN}=f_{AGN}$ [Injected] $-f_{AGN}$ [Zoobot]) of $-0.0018$ and an overall RMSE = 0.013. The solid diagonal line is the 1:1 line. The top plot shows the mean difference and its dispersion as a function of the injected AGN contribution fraction. The colour bar indicates the number of sources in each bin.
@@
Zoobot $0.013\pm 0.005$ $0.076\pm 0.005$ $-0.0018\pm 0.0002$ $0.065\pm 0.004$ $0.061\pm 0.003$
ZoobotM $0.014\pm 0.006$ $0.102\pm 0.009$ $-0.0017\pm 0.0005$ $0.072\pm 0.010$ $0.065\pm 0.010$
ZoobotNM $0.013\pm 0.006$ $0.073\pm 0.005$ $-0.0018\pm 0.0002$ $0.065\pm 0.004$ $0.061\pm 0.004$
GALFIT ($f_{host}$ from original image) $0.12\pm 0.07$ $0.19\pm 0.01$ $-0.018\pm 0.002$ $0.19\pm 0.01$ $0.13\pm 0.01$
GALFITM ($f_{host}$ from original image) $0.13\pm 0.07$ $0.27\pm 0.05$ $-0.024\pm 0.004$ $0.25\pm 0.03$ $0.18\pm 0.02$
GALFITNM ($f_{host}$ from original image) $0.12\pm 0.07$ $0.17\pm 0.01$ $-0.016\pm 0.002$ $0.16\pm 0.01$ $0.11\pm 0.01$
GALFIT ($f_{host}$ from model) $0.17\pm 0.08$ $0.51\pm 0.02$ $-0.075\pm 0.003$ $0.46\pm 0.01$ $0.35\pm 0.01$
GALFITM ($f_{host}$ from model) $0.16\pm 0.09$ $0.48\pm 0.07$ $-0.087\pm 0.007$ $0.54\pm 0.04$ $0.39\pm 0.04$
GALFITNM ($f_{host}$ from model) $0.17\pm 0.08$ $0.52\pm 0.02$ $-0.073\pm 0.003$ $0.45\pm 0.02$ $0.34\pm 0.01$
@@
Here we analyse the performance of the Zoobot model on the subset of 4800 galaxies from the test set. We first calculate the root mean square error (RMSE),
@@
which measures the average difference between the predicted values from a model (e.g. Zoobot or GALFIT) and the actual injected values. In Fig. 6 we show the predicted values of $f_{AGN}$, obtained from Zoobot, versus the real $f_{AGN}$ for all redshifts, which shows a very tight correlation across the whole dynamic range. Quantitatively, we found an overall value of $\text{RMSE}=0.013$ for the Zoobot model, demonstrating very good recovery (both in terms of accuracy and precision) of the injected real AGN contribution fractions. In addition, the top panel of Fig. 6 shows the mean and the dispersion of the difference between the real and predicted values ($\Delta f_{AGN}=f_{AGN}$ [injected] $-f_{AGN}$ [Zoobot]) as a function of the injected AGN fraction, with a value for the mean of $\langle{\Delta f_{AGN}}\rangle=-0.0018$, and dispersion of $\sigma(\Delta f_{AGN})=0.013$. This shows that there is very little bias from the Zoobot predictions to the actual values, and it is independent of $f_{AGN}$, as the difference between the real and predicted value over the whole range of $f_{AGN}$ is always below -0.0043. To check if the performance depends on galaxy structural properties, we further explore whether the Sérsic index impacts how well Zoobot can predict the AGN contribution fraction (see Fig. 15 in Appendix A) and find that for galaxies with $n<1$ the RMSE increases by 30% (compared to the overal RMSE of 0.013), to RMSE$=0.017$ and that high Sérsic indexes ($n>6$) do not lead to worse predictions (i.e., similar RMSE to the overal RMSE of $0.013$).
@@
Figure 7: RMSE as a function of the injected AGN contribution fraction (top left), redshift (top right), S/N (bottom left) and $r_{kron}$ (bottom right). The purple lines correspond to the results from Zoobot and the green lines from GALFIT. The solid lines correspond to the whole sample while the dashed and dotted lines correspond to the mergers and non-merger galaxies, respectively. The error bars show the 95% interval from bootstrapping the RMSE value. The performance of Zoobot is around a factor of 10 better than GALFIT. The shaded region in the top left panel represents the fractional variation (standard deviation divided by the mean) of the PSF, considering the spatial and temporal variations.
@@
Figure 8: Average relative absolute error in bins of injected $f_{AGN}$ (top left panel), redshift (top right panel), S/N (bottom left panel) and $r_{kron}$ (bottom right), from the Zoobot predictions. The error bars correspond to the standard deviation of the data points in each bin. The purple lines correspond to the results from Zoobot and the green, from GALFIT. The solid lines correspond to the whole sample while the dashed and dotted lines correspond to the merger and non-merger galaxies, respectively.